# Randomised benchmarking for universal qudit gates

### v1

@@ 1 Introduction
Driven by the need to exploit the entire Hilbert space available in Physics experiments, the extension qubit information processing to qudits is now experimentally feasible. The main motivation for choosing qudits over qubits is to avoid truncating naturally higher dimensional quantum systems into qubits [1, 2]. Single- and multi-qudit experiments use photons [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], trapped ions [13, 14, 15, 16, 17, 18], superconducting qudits [19, 20, 21, 22, 23, 24], dopants in silicon [25], ultracold atoms [26], and spin systems [27, 28]. Qudits systems are currently in use in quantum communication [29], quantum teleportation [30, 31], quantum memories [32, 33], Bell-state measurements [34], study of spin chains [35, 36, 31, 5, 34, 12], in enhancing quantum error correction techniques [37, 38], in encoding qubits [39] and qudits [40], simulations of many-body systems [41], quantum key distribution [42, 43, 44], simulation of high-energy physics [45, 46, 47, 48, 49, 50, 51, 52, 53], and quantum computing [4, 54, 55]. In particular universal gate sets are necessary to perform quantum computing tasks currently unviable for classical machines. Our work is about benchmarking qudits gates to assess their performance: we extend randomised benchmarking (RB) to estimate the average gate fidelity of single-qudit T gates and controlled-T gates. With our extension of randomised benchmarking to (characterise) T gates, the scalable characterisation of generators of a universal qudit gate set is now available.
@@
Unitary 2-designs are not the only gate sets that can be used to characterise using a randomised benchmarking scheme. Using the techniques of randomised benchmarking, an arbitrary gate set can be characterised [66]. However, such characterisation could in principle require up to $d^{2}$ parameters. This is part of the importance of dihedral benchmarking for qubits [58] and our work for qudits.
@@
A recent scheme that is more general than randomised benchmarking is shadow estimation [68]. For this scheme, real irreducible representations (of a group) are fundamental. Our work, specifically the representation of the rHDG we construct in §3, can also be used—substituting the Clifford gate set—in such a scheme to characterise universal gates.
@@
Our scheme has the following applications. Primarily, it benchmarks single- and multi-qudit T gates. However, our scheme can also be used to characterise any gate of order $d$ diagonal in the computational basis. Additionally, the representation of the rHDG we construct can be exploited in shadow estimation to characterise a universal gate set without requiring an inversion gate. We emphasise that, although our scheme generalises dihedral benchmarking, our techniques in this generalisation are distinct and are of interest by themselves.
@@
Because our work is long and we expect it to be of interest to both experimental and theoretical groups, we explain two roadmaps for our work: one for experimental implementation and the other for theoretical interest. The reader more interested on the implementation of the scheme, §4.2 and §5 contain an illustration of the scheme and the explicit resources (in terms of the number of shots, circuits, and states) required for the implementation of our scheme. In that subsection, we refer to the expressions that are of interest for the characterisation. For a theoretical audience, §3 and, more importantly, §4 are of interest since, in those sections, the gate set is discussed both in creation and application in randomised benchmarking schemes.
@@ 2.1 Setting the scene
In this subsection we introduce notation for the mathematical objects needed for our scheme. Our scheme is developed for $d$-dimensional Hilbert spaces $\mathsf{H}_{d}\coloneqq\mathrm{span}(\ket{i}\colon i\in\operatorname{\mathbb{Z}}_{d})$, with $\operatorname{\mathbb{Z}}_{d}$ integers modulo $d$ [74]. Because currently there is no a clearly defined universal gate set for non-powers of primes, we restrict $d$ to be a power of prime.
Let $\mathscr{M}_{d,d}$ denote the set of $d$-dimensional square matrices with complex entries. Now, let us consider another subset of $\mathscr{M}_{d,d}$: we denote by $\mathscr{D}^{\prime}(d)\coloneqq\{\rho\in\mathscr{M}_{d,d}\colon\operatorname{tr}{\rho}=1,\rho\geq 0,\rho\text{ is hermitian}\}$ the set of density matrices for $d$-dimensional systems.
@@
$\mathscr{B}_{d}\coloneqq\{\mathcal{E}_{\bm{A}}\colon\bm{A}\subset\mathscr{M},\sum_{A\in\bm{A}}A^{\dagger}A=\mathbb{I}_{d}\}.$ (2.3)
When the context is clear, we obviate the operators $\bm{A}$, drop the subindex A, and only write $\mathcal{E}\in\mathscr{B}_{d}$.
@@ 2.2 Pauli and Clifford gates
Some of the basic operations, which can be performed on $\mathsf{H}_{d}$, are represented by the clock and shift matrices ${}_{d}X$ and ${}_{d}Z$ [77, 78]. Let $k$ be any positive integer, and $[k]\coloneqq\{0,1,\ldots,k-1\}$. The matrices ${}_{d}X$ and ${}_{d}Z$ act on the computational basis of a $d$-dimensional Hilbert space as
@@
Another group highly relevant to our work is the Clifford group. We denote the subset of $\mathscr{M}_{d,d}$ formed by unitary matrices by $\mathrm{U}(d)\coloneqq\{u\in\mathscr{M}_{d,d}\colon uu^{\dagger}=\mathbb{I}_{d}\}$. The qudit Clifford group ${}_{d}\mathcal{C}_{2}$ is the normaliser of $\mathcal{P}_{d}$:
@@ 2.3 Universal gates
Universality is firstly a concept defined for a subset of unitary matrices, which is then lifted to a set of gates [80]. Let $\{V_{i}\in\mathrm{U}(d)\}$ be a finite set of unitary matrices. Let $\mathsf{V}\coloneqq\langle V_{i}\rangle$ be the group of matrices generated by the set of matrices $\{V_{i}\}$. Then $\mathsf{V}$ is universal if and only if $\operatorname{closure}(\mathsf{V})\cap\mathrm{U}(d)=\mathrm{U}(d)$ [65, 81], meaning an arbitrary element in $\mathrm{U}(d)$ can be generated to arbitrary precision by some sequence of elements in $\mathsf{V}$.
Each matrix $V\in\mathsf{V}$ induces a channel $\mathcal{E}_{V}\coloneqq V\otimes\bar{V}$. Since $V$ is by construction unitary, we refer to the unitary channel $\mathcal{E}_{V}$ as the representation of the ideal $V$ gate, or as the ideal $V$ gate for short. Similarly, given a set of unitary matrices, such as the Clifford or Pauli groups, we can talk about Clifford or Pauli gates. Therefore, given the clarifications made in the previous two sentences, a universal gate set refers to the physical implementation of a set of matrices $\mathsf{V}$, such that $\mathrm{closure}(\mathsf{V})=\mathrm{U}(d)$.
@@ 2.4 Representation theory detour
There may be a change of basis so the matrices $\mu$ of Eq. (2.4) are made block diagonal; the new basis vectors are thus transformed by the block-diagonal $\mu$ into a subset spanning disjoint subspaces. In such a case, $\mu$ is said to be reducible. If no change of basis can further block diagonalize $\mu$, then each block said to be an irreducible representation of the group. Note that an irreducible block may occur more than once in the reduction of the original representation. More formally: assume there is a list of subspaces $\{\Sigma_{i}\}$ of dimension $\geq 1$ such that $\mu(g)\Sigma_{i}\subset\Sigma_{i}$, v.g. $\mu(g)$ transforms vectors in the block $\Sigma_{i}$ only into combinations of other vectors in $\Sigma_{i}$ and there is no smaller subspace $\Sigma_{i}^{\prime}\subset\Sigma_{i}$ such that $\mu(g)\Sigma_{i}^{\prime}\subset\Sigma_{i}^{\prime}$. Then we can write
@@
We define an important representation in our scheme. For some $d$-dimensional unitary irreducible representation (unirrep) $\gamma$, we define the representation $\mathcal{E}_{\gamma}$ as
@@
In our case, $\gamma(g)\otimes\overline{\gamma(g)}$ will be reducible, as shown below. We also define the noisy version of $\mathcal{E}_{\gamma}$. Let ~$\widetilde{\mathcal{E}}\in\mathscr{B}_{d}$ be as in (2.2). Then the noisy version of $\mathcal{E}_{\gamma}$ is
@@ 2.5 Average gate fidelity and sequence fidelity
Consider a set of channels labelled by the elements of a finite group $\mathbb{G}$. Standard RB schemes estimate the AGF (or related quantities as process fidelity and entanglement fidelity) over a finite group $\mathbb{G}$ [87]. We now recall the formula for the AGF used in gate-independent RB schemes.
@@
For clarity in our exposition we present our results for the case of gate-independent noise, but we notice that our results can be extended to the case where every gate has different noise [90, 91].
@@ 3 Approach
We now introduce the key ingredient of our generalisation of single- and multi-qudit dihedral benchmarking: the qudit real hyperdihedral group (${}_{d}\mathrm{rHDG}$). First we introduce auxiliary groups needed for the definition of the ${}_{d}\mathrm{rHDG}$. Then we construct representations for the auxiliary groups essential to the definition of a unirrep for the ${}_{d}\mathrm{rHDG}$; from this unirrep, our expressions for the gate fidelity and sequence fidelity are derived. This is a good place to make the important observation that the the ${}_{d}\mathrm{rHDG}$’s construction is valid for any diagonal gate. Thus, our scheme is also useful to characterise any diagonal gate with order equal to, or greater than $\operatorname{\mathfrak{o}}(d)$. We conclude this section by presenting the decomposition of the qudit Hilbert space in irreps of ${}_{d}\mathrm{rHDG}$.
@@ 3.1 Constructing the rHDG
The ${}_{d}\mathrm{rHDG}$ requires a product of cyclic groups. The number of cyclic groups in the product and their order depends on $d$ and $T$, with $T$ given in Eq. (2.12). We proceed to describe the structure of an arbitrary finite product of cyclic groups. After introducing the notation, we explain how to construct the specific product for a given diagonal gate $T$.
Some preliminary definitions are necessary before our discussion of the product of cyclic groups. Let $\operatorname{\mathfrak{o}}(d)$ denote the order of order of ${}_{d}T$ in Eq. (2.12); that is,
@@
where—in an abuse of notation—$\alpha_{k}$ is a member of the ring $\operatorname{\mathbb{Z}}_{k}$ [74].
@@
where the symbol $\ltimes$ denotes semidirect product [74, 92]. We use the semidirect product to formally consider the fact that the multiplication rule between elements of the ${}_{d}\mathrm{rHDG}$ is
$S(\sigma_{1})D(\alpha_{1})S(\sigma_{2})D(\alpha_{2})=S(\sigma_{1}\sigma_{2})D(\alpha_{1}^{\sigma_{2}})D(\alpha_{2}),$ (3.8)
noting that $D(\alpha_{1}^{\sigma_{2}})\coloneqq S(\sigma_{2})^{\dagger}D(\alpha_{1})S(\sigma_{2})$ is also a diagonal matrix and is a representative of an element of ×$\bigtimes_{k\in\bm{k}}C_{k}$. In turn, the unirrep $\gamma$ for ${}_{d}\mathrm{rHDG}$ is
@@
We now discuss a fundamental point: the decomposition of the Hilbert space $\mathscr{D}_{d}$ in terms of irreps of ${}_{d}\mathrm{rHDG}$. The representation $\mathcal{E}_{\gamma}=\gamma\otimes\bar{\gamma}$ defined in Eq. (2.15) decomposes the Hilbert space $\mathscr{D}_{d}$ into three irreps; we prove this fact in Appendix C. These representations are: $\Sigma_{\mathbb{I}}$, the trivial irrep; $\Sigma_{0}$, the standard irrep (has dimension $d-1$); and $\Sigma_{+}$, the complement of $\Sigma_{\mathbb{I}}$ and $\Sigma_{0}$. We write the decomposition as
@@
Our scheme also serves to characterise hybrid platforms [93]. A hybrid platform is a combination of two qudits: one is a $d_{1}$-level system and the other a $d_{2}$-level system, with $d_{1}\neq d_{2}$ and also primes or powers of primes. By considering the X and T gates for each subsystem, one of dimension $d_{1}$ and the other of dimension $d_{2}$, we construct the ${}_{d}\mathrm{rHDG}$ for that system. The Kraus-representation of these gates decomposes as in Eq. (3.10). Thus, it is possible to use our scheme to estimate the average gate fidelity of hybrid platforms.
In this section we introduced one irrep for the rHDG that decomposes the Hilbert space into three irreps. Our results are valid not only for a qudit T gate but for any diagonal unitary matrix with order at least $d$. In the following section we use the decomposition in Eq. (3.10) to define our RB scheme. We principally use the projectors onto the irreps in Eq. (3.10).

### v2

@@ 1 Introduction
Driven by the need to exploit the entire Hilbert space available in Physics experiments, the extension from qubit information processing to qudits is now experimentally feasible. The main motivation for choosing qudits over qubits is to avoid truncating naturally higher dimensional quantum systems into qubits [1, 2]. Single- and multi-qudit experiments use photons [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], trapped ions [13, 14, 15, 16, 17, 18], superconducting qudits [19, 20, 21, 22, 23, 24], dopants in silicon [25], ultracold atoms [26], and spin systems [27, 28]. Qudits systems are currently in use in quantum communication [29], quantum teleportation [30, 31], quantum memories [32, 33], Bell-state measurements [34], study of spin chains [35, 36, 31, 5, 34, 12], in enhancing quantum error correction techniques [37, 38], in encoding qubits [39] and qudits [40], simulations of many-body systems [41], quantum key distribution [42, 43, 44], simulation of high-energy physics [45, 46, 47, 48, 49, 50, 51, 52, 53], and quantum computing [4, 54, 55]. In particular universal gate sets are necessary to perform quantum computing tasks currently unviable for classical machines. Our work is about benchmarking qudits gates to assess their performance: we extend randomised benchmarking (RB) to estimate the average gate fidelity of single-qudit T gates and controlled-T gates. With our extension of randomised benchmarking to (characterise) T gates, the scalable characterisation of generators of a universal qudit gate set is now available.
@@
Unitary 2-designs are not the only gate sets that can be used to characterise using a randomised benchmarking scheme. Using the techniques of randomised benchmarking, an arbitrary gate set can be characterised [66]. However, such characterisation could in principle require up to $d^{4}$ parameters. This is part of the importance of dihedral benchmarking for qubits [58] and our work for qudits.
@@
A recent scheme that is more general than randomised benchmarking is shadow estimation [68]. For this scheme, real irreducible representations (of a group) are fundamental. Our work, specifically the representation of the rHDG we construct in §3, can also be used—substituting for the Clifford gate set—in such a scheme to characterise universal gates.
@@
Our scheme has the following applications. Primarily, it benchmarks single- and multi-qudit T gates. However, our scheme can also be used to characterise any gate of order $d$ diagonal in the computational basis satisfying the following condition: in terms of the generators of the diagonal matrices, each diagonal entry parametrised by the generators of the product of cyclic groups, is not a linear combination of the other entries. Additionally, the representation of the rHDG we construct can be exploited in shadow estimation to characterise a universal gate set without requiring an inversion gate. We emphasise that, although our scheme generalises dihedral benchmarking, our techniques in this generalisation are distinct and are of interest by themselves.
@@
Because our work is long and we expect it to be of interest to both experimental and theoretical groups, we explain two roadmaps for our work: one for those readers interested on an experimental implementation and the other for those interested on our theoretical results. The reader more interested on the implementation of the scheme, §4.2 and §5 contain an illustration of the scheme and the explicit resources (in terms of the number of shots, circuits, and states) required for the implementation of our scheme. In that subsection, we refer to the expressions that are of interest for the characterisation. For a theoretical audience, §3 and, more importantly, §4 are of interest since, in those sections, the gate set is discussed both in creation and application in randomised benchmarking schemes.
@@ 2.1 Setting the scene
In this subsection we introduce notation for the mathematical objects needed for our scheme. Our scheme is developed for $d$-dimensional Hilbert spaces $\mathsf{H}_{d}\coloneqq\mathrm{span}(\ket{i}\colon i\in\operatorname{\mathbb{Z}}_{d})$, with $\operatorname{\mathbb{Z}}_{d}$ integers modulo $d$ [74]. Because currently there is no clearly defined universal gate set for non-powers of primes, we restrict $d$ to be a power of prime.
Let $\mathscr{M}_{d,d}$ denote the set of $d$-dimensional square matrices with complex entries. Now, let us consider a subset of $\mathscr{M}_{d,d}$: we denote by $\mathscr{D}^{\prime}(d)\coloneqq\{\rho\in\mathscr{M}_{d,d}\colon\operatorname{tr}{\rho}=1,\rho\geq 0,\rho\text{ is hermitian}\}$ the set of density matrices for $d$-dimensional systems.
@@
$\mathscr{B}_{d}\coloneqq\{\mathcal{E}_{\bm{A}}\colon\bm{A}\subset\mathscr{M}_{d,d},\sum_{A\in\bm{A}}A^{\dagger}A=\mathbb{I}_{d}\}.$ (2.3)
When the context is clear, s we obviate the operators $\bm{A}$, drop the subindex A, and only write $\mathcal{E}\in\mathscr{B}_{d}$.
@@ 2.2 Pauli and Clifford gates
Some of the basic operations that can be performed on $\mathsf{H}_{d}$, are represented by the clock and shift matrices ${}_{d}X$ and ${}_{d}Z$ [77, 78]. Let $k$ be any positive integer, and $[k]\coloneqq\{0,1,\ldots,k-1\}$. The matrices ${}_{d}X$ and ${}_{d}Z$ act on the computational basis of a $d$-dimensional Hilbert space as
@@
Another group highly relevant to our work is the Clifford group. We denote the subset of $\mathscr{M}_{d,d}$ formed by unitary matrices by $\mathrm{U}(d)\coloneqq\{u\in\mathscr{M}_{d,d}\colon uu^{\dagger}=u^{\dagger}u=\mathbb{I}_{d}\}$. The qudit Clifford group ${}_{d}\mathcal{C}_{2}$ is the normaliser of $\mathcal{P}_{d}$:
@@ 2.3 Universal gates
Universality is firstly a concept defined for a subset of unitary matrices, which is then lifted to a set of gates [80]. Let $\{V_{i}\in\mathrm{U}(d)\}$ be a finite set of unitary matrices. Let $\mathsf{V}\coloneqq\langle V_{i}\rangle$ be the group of matrices generated by the set of matrices $\{V_{i}\}$. Then $\mathsf{V}$ is universal if and only if $\operatorname{closure}(\mathsf{V})\cap\mathrm{U}(d)=\mathrm{U}(d)$ [65, 81], meaning an arbitrary element in $\mathrm{U}(d)$ can be generated to arbitrary precision by some finite sequence of elements in $\mathsf{V}$.
Each matrix $V\in\mathsf{V}$ induces a channel $\mathcal{E}_{V}\coloneqq V\otimes\bar{V}$. Since $V$ is by construction unitary, we refer to the unitary channel $\mathcal{E}_{V}$ as the representation of the ideal $V$ gate, or as the ideal $V$ gate for short. In the same manner, given a set of unitary matrices, such as the Clifford or Pauli groups, we can talk about Clifford or Pauli gates. Therefore, given the clarifications made in the previous two sentences, a universal gate set refers to the physical implementation of a set of matrices $\mathsf{V}$, such that $\mathrm{closure}(\mathsf{V})=\mathrm{U}(d)$.
@@ 2.4 Representation theory detour
There may be a change of basis so the matrices $\mu$ of Eq. (2.4) are made block diagonal; the new basis vectors are thus transformed by the block-diagonal $\mu$ into a subset spanning disjoint subspaces. In such a case, $\mu$ is said to be reducible. If no change of basis can further block-diagonalize $\mu$, then each block said to be an irreducible representation of the group. Note that an irreducible block may occur more than once in the reduction of the original representation. More formally: assume there is a list of subspaces $\{\Sigma_{i}\}$ of dimension $\geq 1$ such that $\mu(g)\Sigma_{i}\subset\Sigma_{i}$, v.g. $\mu(g)$ transforms vectors in the block $\Sigma_{i}$ only into combinations of other vectors in $\Sigma_{i}$ and there is no smaller subspace $\Sigma_{i}^{\prime}\subset\Sigma_{i}$ such that $\mu(g)\Sigma_{i}^{\prime}\subset\Sigma_{i}^{\prime}$. Then we can write
@@
We define an important representation in our scheme. For some $d$-dimensional unitary irreducible representation (unirrep) $\gamma$ of $\mathbb{G}$, we define the representation $\mathcal{E}_{\gamma}$ as
@@
In our case, $\gamma(g)\otimes\overline{\gamma(g)}$ will be reducible, as shown below. We also define the noisy representative of $\mathcal{E}_{\gamma}$. Let ~$\widetilde{\mathcal{E}}\in\mathscr{B}_{d}$ be as in (2.2). Then the noisy representative of $\mathcal{E}_{\gamma}$ is
@@ 2.5 Average gate fidelity and sequence fidelity
Consider a set of channels labelled by the elements of a finite group $\mathbb{G}$. Standard RB schemes estimate the AGF (or related quantities such as process fidelity and entanglement fidelity) over $\mathbb{G}$ [87]. We now recall the formula for the AGF used in gate-independent RB schemes.
@@
For clarity in our exposition we present our results for the case of gate-independent noise, but we note that our results can be extended to the case where every gate has different noise [90, 91].
@@ 3 Approach
We now introduce the key ingredient of our generalisation of single- and multi-qudit dihedral benchmarking: the qudit real hyperdihedral group (${}_{d}\mathrm{rHDG}$). First we introduce auxiliary groups needed for the definition of the ${}_{d}\mathrm{rHDG}$. Then we construct representations for the auxiliary groups essential to the definition of a unirrep for the ${}_{d}\mathrm{rHDG}$; from this unirrep, our expressions for the gate fidelity and sequence fidelity are derived. This is a good place to make the important observation that the ${}_{d}\mathrm{rHDG}$’s construction is valid for any diagonal gate. Thus, our scheme is also useful to characterise any diagonal gate with order equal to, or greater than two, such that its parametrisation does not have linearly independent entries. Alternatively, any diagonal matrix with order equal to, or greater than $\operatorname{\mathfrak{o}}(d)$ as defined in Eq. (3.1) can also be characterised. We conclude this section by presenting the decomposition of the qudit Hilbert space in irreps of ${}_{d}\mathrm{rHDG}$.
@@ 3.1 Constructing the rHDG
The ${}_{d}\mathrm{rHDG}$ requires a product of cyclic groups. The number of cyclic groups in the product and their order depends on $d$ and ${}_{d}T$, with ${}_{d}T$ given in Eq. (2.12). We proceed to describe the structure of an arbitrary finite product of cyclic groups. After introducing the notation, we explain how to construct the specific product for a given diagonal gate $T$.
Some preliminary definitions are necessary before our discussion of the product of cyclic groups. Let $\operatorname{\mathfrak{o}}(d)$ denote the order of ${}_{d}T$ in Eq. (2.12); that is,
@@
where—in an abuse of notation—$\alpha_{k}$ can also denote a member of the ring $\operatorname{\mathbb{Z}}_{k}$ [74].
@@
where the symbol $\ltimes$ denotes semidirect product [74, 92]. We use the semidirect product to formally consider the multiplication rule between elements of the ${}_{d}\mathrm{rHDG}$ is
$\operatorname{\mathsf{S}}(\sigma_{1})\operatorname{\mathsf{D}}(\alpha_{1})\operatorname{\mathsf{S}}(\sigma_{2})\operatorname{\mathsf{D}}(\alpha_{2})=\operatorname{\mathsf{S}}(\sigma_{1}\sigma_{2})\operatorname{\mathsf{D}}(\alpha_{1}^{\sigma_{2}})\operatorname{\mathsf{D}}(\alpha_{2}),$ (3.8)
noting that $\operatorname{\mathsf{D}}(\alpha_{1}^{\sigma_{2}})\coloneqq\operatorname{\mathsf{S}}(\sigma_{2})^{\dagger}\operatorname{\mathsf{D}}(\alpha_{1})\operatorname{\mathsf{S}}(\sigma_{2})$ is also a diagonal matrix and is a representative of an element in ×$\bigtimes_{k\in\bm{k}}C_{k}$. In turn, the unirrep $\gamma$ for ${}_{d}\mathrm{rHDG}$ is
@@
We now discuss a fundamental point: the decomposition of the Hilbert space $\mathscr{D}_{d}$ in terms of irreps of ${}_{d}\mathrm{rHDG}$. The representation $\mathcal{E}_{\gamma}=\gamma\otimes\bar{\gamma}$ defined in Eq. (2.15) decomposes the Hilbert space $\mathscr{D}_{d}$ into three irreps; we prove this fact in Appendix C. These representations are: $\Sigma_{\mathbb{I}}$, the trivial irrep; $\Sigma_{0}$, the standard irrep (of dimension $d-1$); and $\Sigma_{+}$, the complement of $\Sigma_{\mathbb{I}}$ and $\Sigma_{0}$. We write the decomposition as
@@
Our scheme also serves to characterise hybrid platforms [93]. A hybrid platform is a combination of two qudits: one is a $d_{1}$-level system and the other a $d_{2}$-level system, with $d_{1}\neq d_{2}$ and also primes or powers of primes. By considering the X and T gates for each subsystem, one of dimension $d_{1}$ and the other of dimension $d_{2}$, we construct the ${}_{d}\mathrm{rHDG}$ for the combined system. The Kraus-representation of these gates decomposes as in Eq. (3.10). Thus, it is possible to use our scheme to estimate the average gate fidelity of hybrid platforms.
In this section we introduced one irrep for the rHDG that leads, by using the associated Pauli-Liouville representation, to the decomposition of the Hilbert space $\mathscr{D}_{d}$ into three irreps. Our results are valid not only for a qudit T gate but for any diagonal unitary matrix with order at least 2, such that the parametrisation of a representative of the product of cyclic groups has no linearly independent entries. In the following section we use the decomposition in Eq. (3.10) to define our RB scheme. We principally use the projectors onto the irreps in Eq. (3.10).